# Search for transient gamma-ray emission from magnetar flares using Fermi-LAT

### v1

@@ III.2 Results from other magnetars
Some of the magnetars are closely associated with Fermi-LAT 4FGL-DR3 point sources, which are in near proximity to the magnetars. These include Swift J1834.9-0846, 1E 1547.0-5408, SGR 1745-2900, 1E 2259+586 and SGR 1833-0832. These are shown using a white-plus marker, while the X-ray positions of the magnetars are depicted using a red diamond. The spectral spatial model used in the analyses of these sources also accounts for the -ray emission from these Fermi-LAT 4FGL-DR3 point sources. The closest Fermi-DR3 point sources are 4FGL J1834.5-0846e (spp)444“spp” stands for “SNR (Supernova remnant), Pulsar or PWN (Pulsar wind nebula)” and refers to sources of unknown nature but overlapping with known SNRs or PWNe and thus candidate members of these classes Abdollahi et al. (2022), associated with Swift J1834; 4FGL J1550.8-5424c (spp), associated with 1E 1547; 4FGL J1745.6-2859 (Galactic center), associated with SGR 1745-2900; 4FGL J2301.9+5855e (Supernova remnant), associated with 1E 2259; and 4FGL J1834.2-0827c*, associated with SGR 1833. The source classes of the nearest sources are given in parentheses, while the unclassified sources are denoted by an asterisk. More details about these nearby 4FGL-DR3 point sources and source classes can be found in Abdollahi et al. (2022). Apart from these known sources, no other statistically significant -ray signal with significance is detected within a one month span of the X-ray flares for each of these magnetars. The 95% confidence level energy flux limits (0.1 - 300 GeV), for all of the magnetars can be found in Table 1.
@@ IV Conclusions
In this work, we have searched for transient gamma-ray emission between 0.1-300 GeV from magnetar flares using 12 years of Fermi-LAT data. For this purpose, we searched in temporal windows of 1 day and 15 days around 15 distinct flares from 11 different magnetars. The times of magnetar outbursts were obtained using X-ray data based on the compilation in Coti Zelati et al. (2018). We do not detect any gamma-ray emission above in coincidence with almost all the flares. The corresponding 95% upper limits on the energy flux for these can be found in Table 1 and are The only magnetar for which we see gamma-ray emission above is for 1E 1048.1-5937, for which the gamma-ray signal is seen around 10 days after the peak of the X-ray flare with observed energy flux of around and , in the pertinent time bins. The gamma-ray signal is spread over 4.13 days and corresponds to two distinct flares that last about seven hours each, separated by 1.17 days. However, given the proximity of this source to the galactic plane, the signal could be caused by contamination from diffuse emission.
@@ Acknowledgements.
The authors thank Siddhant Manna for useful discussions.References
@@
- Abdo et al. (2010) A. A. Abdo, M. Ackermann, M. Ajello, A. Allafort, L. Baldini, J. Ballet, G. Barbiellini, M. G. Baring, D. Bastieri, R. Bellazzini, et al., Astrophys. J. Lett. 725, L73 (2010), eprint 1011.0091.
@@
- Wilks (1938) S. S. Wilks, The annals of mathematical statistics 9, 60 (1938).

### v2

@@ III.2 Results from other magnetars
Some of the magnetars are close to Fermi-LAT 4FGL-DR3 point sources, which are in near proximity to the magnetars. These include Swift J1834.9-0846, 1E 1547.0-5408, SGR 1745-2900, 1E 2259+586 and SGR 1833-0832. These are shown using a white-plus marker, while the X-ray positions of the magnetars are depicted using a red diamond. The spectral spatial model used in the analyses of these sources also accounts for the -ray emission from these Fermi-LAT 4FGL-DR3 point sources. The closest Fermi-DR3 point sources are 4FGL J1834.5-0846e (spp)444“spp” stands for “SNR (Supernova remnant), Pulsar or PWN (Pulsar wind nebula)”, and refers to sources of unknown nature but overlapping with known SNRs or PWNe and thus candidate members of these classes Abdollahi et al. (2022), associated with Swift J1834; 4FGL J1550.8-5424c (spp), associated with 1E 1547; 4FGL J1745.6-2859 (Galactic center), associated with SGR 1745-2900; 4FGL J2301.9+5855e (Supernova remnant), associated with 1E 2259; and 4FGL J1834.2-0827c*, associated with SGR 1833. The source classes of the nearest sources are given in parentheses, while the unclassified sources are denoted by an asterisk. More details about these nearby 4FGL-DR3 point sources and source classes can be found in Abdollahi et al. (2022). Apart from these known sources, no other statistically significant -ray signal with significance is detected within a one month span of the X-ray flares for each of these magnetars. The 95% confidence level energy flux limits (0.1 - 10 GeV), for all of the magnetars can be found in Table 1. Finally, for some of the magnetars for which the X-ray flares last for greater than 600 days, we also did an extended search for a gamma-ray signal in a +200 days window starting from the peak of the flare. However, we do not detect any signal (with ) in such an extended window.
@@ IV Conclusions
In this work, we have searched for transient gamma-ray emission between 0.1-10 GeV from magnetar flares using 12 years of Fermi-LAT data. For this purpose, we searched in three temporal windows of 1 day, 7 days, and 15 days around 15 distinct flares from 11 different magnetars. The times of magnetar outbursts were obtained using X-ray data based on the compilation in Coti Zelati et al. (2018). We do not detect any gamma-ray emission above in coincidence with almost all the flares. The corresponding 95% upper limits on the energy flux for these can be found in Table 1 and are between . We note that the upper limits are of the same order of magnitude as those obtained in Abdo et al. (2010b); Li et al. (2017). The only magnetar for which we see gamma-ray emission (with significance of 4.4) is for 1E 1048.1-5937, for which the gamma-ray signal is seen around 10 days after the peak of the X-ray flare with observed energy flux of around and , in the pertinent time bins. However, given the proximity of this source to the galactic plane, the signal could be caused by contamination from diffuse emission. Additionally, the presence of the X-ray binary 2E-2336 within the 99% uncertainty region suggests that it could be a potential contributor to the detected gamma-ray signal.
Therefore, except for one object, we can definitely rule out gamma-ray emission in coincidence with the magnetar X-ray outbursts with our analysis. As discussed in Abdo et al. (2010b) and Li et al. (2017), the non-detections rule out most of the parameter space of the outer gap model for magnetars proposed in Cheng and Zhang (2001). Our flux limits for 1E2259+586 are of the same order and magnitude as Li et al. (2017) and also rule out the model for GeV gamma-ray emission proposed in Takata et al. (2013).
In a forthcoming work, we shall extend this search to extragalactic magnetars around the time of giant flares, such as GRB 200415A and GRB 231115A.
@@ Acknowledgements.
The authors thank Siddhant Manna for useful discussions and Raniere Menezes for help with easyFermi. We are also grateful to the anonymous referees for very useful feedback on the manuscript.References
@@
- Rea and De Grandis (2025) N. Rea and D. De Grandis, arXiv e-prints arXiv:2503.04442 (2025), eprint 2503.04442.
@@
- Abdo et al. (2010a) A. A. Abdo, M. Ackermann, M. Ajello, A. Allafort, L. Baldini, J. Ballet, G. Barbiellini, M. G. Baring, D. Bastieri, R. Bellazzini, et al., Astrophys. J. Lett. 725, L73 (2010a), eprint 1011.0091.
@@
- Abdo et al. (2010b) A. A. Abdo, M. Ackermann, M. Ajello, A. Allafort, E. Antolini, W. B. Atwood, M. Axelsson, L. Baldini, J. Ballet, G. Barbiellini, et al., Astrophys. J. Supp. 188, 405 (2010b), eprint 1002.2280.
@@
- Ballet et al. (2023) J. Ballet, P. Bruel, T. H. Burnett, B. Lott, and The Fermi-LAT collaboration, arXiv e-prints arXiv:2307.12546 (2023), eprint 2307.12546.